# A NICER View of PSR J1231--1411: A Complex Case

### v1

@@ Abstract
Recent constraints on neutron star mass and radius have advanced our understanding of the equation of state of cold dense matter. Some of them have been obtained by modeling the pulses of three millisecond X-ray pulsars observed by the Neutron Star Interior Composition Explorer (NICER). Here, we present a Bayesian parameter inference for a fourth pulsar, PSR J1231--1411, using the same technique with NICER and XMM-Newton data. When applying a broad mass-inclination prior from radio timing measurements and the emission region geometry model that can best explain the data, we find likely-converged results only when using a limited radius prior. If limiting the radius to be consistent with the previous observational constraints and equation of state analyses, we infer the radius to be $12.6\pm 0.3$ km and the mass to be $1.04_{-0.03}^{+0.05}$ $M_{\odot}$, each reported as the posterior credible interval bounded by the $16\,\%$ and $84\,\%$ quantiles. If using an uninformative prior but limited between $10$ and $14$ km, we find otherwise similar results, but $R_{\mathrm{eq}}=13.5_{-0.5}^{+0.3}$ km for the radius. In both cases, we find a non-antipodal hot region geometry where one emitting spot is at the equator or slightly above, surrounded by a large colder region, and where a non-circular hot region lies close to southern rotational pole. If using a wider radius prior, we only find solutions that fit the data significantly worse. We discuss the challenges in finding the better fitting solutions, possibly related to the weak interpulse feature in the pulse profile.
journal: The Astrophysical Journalfacilities: NICER, XMMsoftware: Cython (Behnel et al., 2011), fgivenx (Handley, 2018), GetDist (Lewis, 2019), GNU Scientific Library (GSL; Gough 2009), HEASoft (Nasa High Energy Astrophysics Science Archive Research Center (2014), Heasarc), Matplotlib (Hunter, 2007), MPI for Python (Dalcín et al., 2008), MultiNest (Feroz et al., 2009), nestcheck (Higson, 2018), NumPy (van der Walt et al., 2011),PyMultiNest (Buchner et al., 2014), Python/C language (Oliphant, 2007), SciPy (Jones et al., 2001), X-PSI (Riley et al., 2023).
@@ 1 Introduction
Modeling the X-ray pulses from a neutron star (NS) can be used to infer the NS mass, radius, and the geometric properties of the X-ray emitting regions (see e.g., Pavlov & Zavlin, 1997; Watts et al., 2016; Bogdanov et al., 2019b, and the references therein). The inferred mass and radius can then be used to constrain the equation of state (EOS) of the supranuclear dense matter in neutron star cores (see e.g., Lattimer & Prakash, 2007). On the other hand, the emitting regions are associated with the NS polar caps, and their properties can be used to probe the NS magnetic field structure (e.g., Bilous et al., 2019; Carrasco et al., 2023).
Previously, X-ray pulses of three NSs have been modeled using Neutron Star Interior Composition Explorer (NICER; Gendreau et al., 2016) observations: PSR J0030$+$0451 (Miller et al., 2019; Riley et al., 2019; Salmi et al., 2023; Vinciguerra et al., 2024), the massive pulsar PSR J0740$+$6620 (Miller et al., 2021; Riley et al., 2021; Salmi et al., 2022, 2023; Dittmann et al., 2024; Salmi et al., 2024), and the bright pulsar PSR J0437--4715 (Choudhury et al., 2024a). These have already provided useful constraints on dense matter models (see e.g., Miller et al., 2021; Raaijmakers et al., 2021; Biswas, 2022; Annala et al., 2023; Han et al., 2023; Takátsy et al., 2023; Rutherford et al., 2024). However, measuring masses and radii for a larger set of NSs and over a broad range of masses is expected to allow both tighter and more robust constraints.
PSR J1231--1411 is a rotation-powered millisecond pulsar (RMP, discovered by Ransom et al. 2011), like many NSs previously analyzed with NICER. Its thermally emitting surface regions are expected to be heated by the bombardment of charged particles from a magnetospheric return current (see, e.g., Ruderman & Sutherland, 1975; Arons, 1981; Harding & Muslimov, 2001). Soft X-ray pulsations at the known spin frequency of $271$ Hz were found by Ray et al. (2019) and further analyzed in Bogdanov et al. (2019a). The pulsed emission shows a broad and slightly asymmetric main pulse and a much fainter secondary interpulse (see Figure 1). This feature differs from those seen in NICER RMPs analyzed so far (see Bogdanov et al. 2019a for PSR J0030$+$0451 and PSR J0437--4715, and Wolff et al. 2021 for PSR J0740$+$6620). Similar to PSR J0740$+$6620 and PSR J0437--4715, PSR J1231--1411 is located in a binary system, allowing radio timing measurements to constrain the NS mass and observer inclination based on the relativistic Shapiro delay (Fonseca et al., 2021; Reardon et al., 2024). However, the obtained constraints in the preliminary radio timing results (available at the time of this work) are far less restrictive for PSR J1231--1411 (Cromartie et al., 2024) compared to PSR J0740$+$6620 and PSR J0437--4715. In this work, we use these measurements as prior distributions and explore the robustness of NS parameter inference for PSR J1231--1411.
@@
Figure 1: The phase-folded PSR J1231--1411 event data for two rotational cycles (for clarity). The top panel shows the pulse profile summed over the channels. The total number of counts is given by the sum over all phase-channel pairs (over both cycles). For the modeling all the event data are grouped into a single rotational cycle instead, and thus each phase-channel bin has twice the number of counts shown here.
Figure 2: XMM-Newton EPIC-pn camera PSR J1231--1411 count number spectrum as a function of nominal photon energies (blue stepped curve). The measured rescaled background (red stepped curve) and the background limits used in the analysis (purple dashed and dot-dashed stepped curves) are also shown. The choice of the limits is explained in Section 3.4. The complete figure set (3 images), including also the spectra and backgrounds for EPIC-MOS1 and EPIC-MOS2 instruments is available in the online journal (HTML version).
@@ 2.1 NICER
The NICER dataset for this work was produced using HEASoft 6.32.1111https://heasarc.gsfc.nasa.gov/ftools (Nasa High Energy Astrophysics Science Archive Research Center (2014), Heasarc) and CALDB xti20221001. We selected data from the full NICER mission up through 2023 May 14 (ObsID 6060060730), which is before NICER was affected by the light leak [^2] https://heasarc.gsfc.nasa.gov/docs/nicer/analysis_threads/light-leak-overview/. After standard L2 processing with the task nicerl2, but before any of our filtering, the initial dataset contained about 3.53 Ms of exposure time. We applied further cuts to exclude data with magnetic cutoff rigidity $<1.5$ GeV/$c$; planetary K-index $K_{p}>5$; overshoot rate [^3] https://heasarc.gsfc.nasa.gov/docs/nicer/analysis_threads/overshoot-intro/ $>1.5$ counts per second (cps) per Focal Plane Module (FPM); or a median undershoot rate [^4] https://heasarc.gsfc.nasa.gov/docs/nicer/analysis_threads/undershoot-intro/ $>200$ cps/FPM. These cuts reduce the exposure to 81% of the L2 exposure. Subsequently, we generated a background lightcurve using only 2–10 keV photons (where the pulsar flux is small so the count rate is background dominated) with 16-second bins and discarded times where that rate exceeded 1.25 cps. This removes the times polluted by high levels of trapped electron (TREL, in the nomenclature of the SCORPEON background model, see below) and low energy electron (LEEL) backgrounds, reducing the exposure to 75% of the L2 exposure. The final exposure was 2651393.0413 seconds.
@@
The timing model to assign pulse phases to photon measured times was developed by first fitting a timing model to the radio time of arrivals from the Nançay radio telescope (NRT). Details on the NRT and on pulsar observations with this instrument can be found in Guillemot et al. (2023). This determined the astrometric and binary parameters (including Shapiro delay) and a constant dispersion measure. These were then held fixed while the frequency and frequency first derivative were fit to the Fermi LAT gamma ray data, as done in FERMI-LAT Collaboration et al. (2022). Finally, the reference epoch was set to the values used in the Third Fermi LAT Pulsar Catalog (3PC; Smith et al. 2023) so the absolute pulse phases will be consistent with that. This model provides pulse phases accurate to $<0.1\%$ of a pulse period over the entire duration of the NICER dataset.
@@
The NICER science tool nicerl3-spect was used to compute a pulse height spectrum, response matrix, and background estimate using the SCORPEON background model [^5] https://heasarc.gsfc.nasa.gov/docs/nicer/analysis_threads/scorpeon-overview/. We used XSPEC (Arnaud, 1996) to fit the SCORPEON model parameters simultaneously with different source spectral models, including two blackbody components, one blackbody and one powerlaw component (with a powerlaw index fixed at $1.8$ as in Ransom et al. 2011), just a powerlaw component, one neutron star atmosphere component (nsatmos), and two neutron star atmosphere components. These all were modified by interstellar absorption (e.g., TBabs(BB + powerlaw)). For these data, we found that, in addition to the standard free parameters in the SCORPEON model, we needed to free the normalization of the Ovii line at 0.57 keV (presumably originating from solar wind charge exchange). In prior work before the development of SCORPEON (Ray et al., 2019), this line was added as an ad hoc Gaussian line. We found that the obtained SCORPEON background estimates were notably dependent on the chosen source model, and therefore we predicted the background to range from the smallest estimate to the highest at each channel (after including also $1$ standard deviation uncertainty for each model). This range was used only in the end to compare our inferred background to it (see Section 4.2).
@@
In the pulse profile analysis, we used the pulse invariant (PI) channel subset [30,150), corresponding to the nominal photon energy range [0.3, 1.5] keV, and 32 rotational phase bins. The data split over 2 rotational cycles are visualized in Figure 1.
@@ 2.2 XMM-Newton
The XMM-Newton EPIC-MOS (Turner et al. 2001; 1 and 2) and EPIC-pn (Strüder et al., 2001) data (ObsID 0605470201) used in this study were obtained on 2009 July 15, as part of a Fermi/LAT unidentified source observation campaign (Ransom et al., 2011). We utilize the Scientific Analysis System software (SAS) [^6] https://www.cosmos.esa.int/web/xmm-newton/sas to analyze and build our data products that in turn are input into the inference software. The fields of the new unidentified LAT sources were observed using the XMM-Newton imaging mode, thus with insufficient time resolution to look for pulsations. Using the original raw data products from the EPIC-MOS and EPIC-pn observations we extracted X-ray events from the now-known pulsar position. We set the PATTERN $\leq$ 12 for the EPIC-MOS1 and EPIC-MOS2 instruments and PATTERN $\leq$ 4 for the EPIC-pn along with FLAG=0 for all three instruments. This resulted in 28.8 ks of exposure from the EPIC-MOS instruments and 17.7 ks of exposure for the EPIC-pn instrument. For the EPIC-MOS1 and EPIC-MOS2 event data we chose source+background extraction regions of 32 arc seconds and source-free background extraction regions of 90 arc seconds. The extraction regions for source+background and background events for the EPIC-pn were 38 and 60 arc seconds, respectively. We filtered out times of higher than average particle noise from each event stream. The resulting final list of events are stored in FITS files that are then converted into ASCII files containing just the event times and pulse invariant (PI) value. However, no pulse phases are included in the XMM-Newton event files because the observations were made with a detector frame time too long to detect pulsations. Once the event filtering is selected we build X-ray spectra using the SAS tool “evselect” and then use the SAS tools “rmfgen” and “arfgen” to build response matrices. The resulting response matrices are then turned into ASCII text files for ingesting into the inference code that is described below. We also implicitly make the assumption that the average pulsar X-ray properties have not changed from the epoch of the XMM-Newton observation to the epoch of the NICER observations. The assumption is justified since no evidence for long-term flux variability of PSR J1231--1411 was found in Bogdanov et al. (2019a, see Section 7.3 there).
@@ 3 Modeling Procedure
We use the X-ray Pulse Simulation and Inference [^7] https://github.com/xpsi-group/xpsi (X-PSI) code, with version v2.1.1 for the inference runs and v2.2.7 for producing the figures (Riley et al., 2023). Complete information of each run, data products, posterior sample files, and all of the analysis files can be found in the Zenodo repository (Salmi et al., 2024). In the next sections we summarize the modeling procedure and focus on how it differs from that used in previous work.
@@ 3.1 Effective-area Scaling Models
As in previous works (Riley et al., 2021; Choudhury et al., 2024a; Salmi et al., 2024; Vinciguerra et al., 2024), we capture the uncertainty in the effective area of both NICER and all XMM-Newton detectors by defining absolute energy-independent effective-area scaling factors. Each factor consists of a shared and telescope specific component (see Section 2.2. in Salmi et al. 2024), and as in most recent works, we select the uncertainties in those to be 10% and 3%, correspondingly, leading to a 10.4% uncertainty in the full scaling factor. As before, we commonly assume all scaling factors to be identical between EPIC-pn, EPIC-MOS1 and EPIC-MOS2 (leaving then only the scaling factors $\alpha_{\mathrm{NICER}}$ and $\alpha_{\mathrm{XMM}}$ as free parameters). Since this may not be true, we have also performed a test run treating $\alpha_{\mathrm{XMM}}$ with $\alpha_{\mathrm{pn}}$, $\alpha_{\mathrm{MOS1}}$, and $\alpha_{\mathrm{MOS2}}$ as free independent parameters, although with all of them defined to have the same uncertainty (10.4% as mentioned above) and the same correlation with $\alpha_{\mathrm{NICER}}$ (see Section 4.1).
@@ 3.2 Pulse Profile Modeling Using X-PSI
As in previous NICER analyses, we use the ‘Oblate Schwarzschild’ approximation for the space-time and NS shape model to calculate the energy-resolved X-ray pulses (see e.g., Miller & Lamb, 1998; Nath et al., 2002; Poutanen & Gierliński, 2003; Cadeau et al., 2007; Morsink et al., 2007; Lo et al., 2013; AlGendy & Morsink, 2014; Bogdanov et al., 2019b; Watts, 2019). The spectrum and the beaming pattern of the surface radiation are in most cases assumed to follow those from a fully-ionized hydrogen atmosphere model NSX (Ho & Lai, 2001), but we also performed a run assuming fully-ionized helium NSX atmosphere (see Section 4.2). In addition, for one case we also modeled the NS surface outside the hot regions, assuming it emits blackbody radiation (for computational efficiency). This assumption is more realistic than having no surface emission, although more accurate still would be to use a numerical (and thus more expensive) partially-ionized atmosphere model for the likely relatively cold surface. To account for the interstellar attenuation, we use the TBabs model (Wilms et al., 2000, updated in 2016).
To characterize the geometry (and number) of the hot regions, we apply two different models: Single Temperature – Unshared (ST-U) and Protruding Dual Temperature – Unshared (PDT-U). In the former model two hot regions are described by two single-temperature disjoint spherical caps with unshared parameters. The latter model is the most complex model applied with X-PSI so far, where each of the two hot regions is additionally described by two overlapping spherical caps with different temperatures (see Riley et al. 2019 and Figure 1 of Vinciguerra et al. 2023 for more details). Where the two overlap, emission of only one of them is accounted for, and that component is referred to as being “superseding” (the other being “ceding”).
For the pulse profile modeling resolution with X-PSI we mostly follow the settings [^8] For more details see Section 2.3.1 in Vinciguerra et al. (2024) or X-PSI documentation at https://xpsi-group.github.io/xpsi/hotregion.html. used in the previous PSR J0740$+$6620 analyses (Riley et al., 2021; Salmi et al., 2022, 2024). For example, we calculate multiple images (photons reaching the observer from the same point at the NS surface with many deflection angles due to high gravity) up to third order in case they are visible. The initial deflection angle grid is again calculated for 512 different emission angles at each co-latitude. However, to speed-up the calculation we reduced the number of model energies (from 128 to 64) and phases (from 64 to 32) for the more expensive PDT-U model. This change was tested and found to have no effects on the results for a ST-U run. In contrast, the resolution of the hot region cell mesh (sqrt_num_cells = 32, min_sqrt_num_cells = 10, max_sqrt_num_cells = 64) was not reduced, to avoid loss of accuracy in the computed waveforms for extreme configurations as found by Choudhury et al. (2024b).

### v2

@@ Abstract
Recent constraints on neutron star mass and radius have advanced our understanding of the equation of state (EOS) of cold dense matter. Some of them have been obtained by modeling the pulses of three millisecond X-ray pulsars observed by the Neutron Star Interior Composition Explorer (NICER). Here, we present a Bayesian parameter inference for a fourth pulsar, PSR J1231--1411, using the same technique with NICER and XMM-Newton data. When applying a broad mass-inclination prior from radio timing measurements and the emission region geometry model that can best explain the data, we find likely converged results only when using a limited radius prior. If limiting the radius to be consistent with the previous observational constraints and EOS analyses, we infer the radius to be $12.6\pm 0.3$ km and the mass to be $1.04_{-0.03}^{+0.05}$ $M_{\odot}$, each reported as the posterior credible interval bounded by the $16\,\%$ and $84\,\%$ quantiles. If using an uninformative prior but limited between $10$ and $14$ km, we find otherwise similar results, but $R_{\mathrm{eq}}=13.5_{-0.5}^{+0.3}$ km for the radius. In both cases, we find a nonantipodal hot region geometry where one emitting spot is at the equator or slightly above, surrounded by a large colder region, and where a noncircular hot region lies close to southern rotational pole. If using a wider radius prior, we only find solutions that fit the data significantly worse. We discuss the challenges in finding the better fitting solutions, possibly related to the weak interpulse feature in the pulse profile.
journal: The Astrophysical Journalfacilities: NICER, XMMsoftware: Cython (Behnel et al., 2011), fgivenx (Handley, 2018), GetDist (Lewis, 2019), GNU Scientific Library (GSL; Galassi et al. 2009), HEASoft (Nasa High Energy Astrophysics Science Archive Research Center (2014), Heasarc), Matplotlib (Hunter, 2007), MPI for Python (Dalcín et al., 2008), MultiNest (Feroz et al., 2009), nestcheck (Higson, 2018), NumPy (van der Walt et al., 2011),PyMultiNest (Buchner et al., 2014), Python/C language (Oliphant, 2007), SciPy (Jones et al., 2001), X-PSI (Riley et al., 2023).
@@ 1 Introduction
Modeling the X-ray pulses from a neutron star (NS) can be used to infer the NS’s mass, radius, and the geometric properties of the X-ray emitting regions (see e.g., Pavlov & Zavlin, 1997; Watts et al., 2016; Bogdanov et al., 2019b, and the references therein). The inferred mass and radius can then be used to constrain the equation of state (EOS) of the supranuclear dense matter in NS cores (see e.g., Lattimer & Prakash, 2007). On the other hand, the emitting regions are associated with the NS’s polar caps, and their properties can be used to probe the NS’s magnetic field structure (e.g., Bilous et al., 2019; Carrasco et al., 2023).
Previously, X-ray pulses of three NSs have been modeled using Neutron Star Interior Composition Explorer (NICER; Gendreau et al., 2016) observations: PSR J0030$+$0451 (Miller et al., 2019; Riley et al., 2019; Salmi et al., 2023; Vinciguerra et al., 2024), the massive pulsar PSR J0740$+$6620 (Miller et al., 2021; Riley et al., 2021; Salmi et al., 2022, 2023; Dittmann et al., 2024; Salmi et al., 2024), and the bright pulsar PSR J0437--4715 (Choudhury et al., 2024a). These have already provided useful constraints on dense matter models (see, e.g., Miller et al., 2021; Raaijmakers et al., 2021; Biswas, 2022; Annala et al., 2023; Han et al., 2023; Takátsy et al., 2023; Rutherford et al., 2024). However, measuring masses and radii for a larger set of NSs and over a broad range of masses is expected to allow both tighter and more robust constraints.
PSR J1231--1411 is a rotation-powered millisecond pulsar (RMP; discovered by Ransom et al. 2011), like many NSs previously analyzed with NICER. Its thermally emitting surface regions are expected to be heated by the bombardment of charged particles from a magnetospheric return current (see, e.g., Ruderman & Sutherland, 1975; Arons, 1981; Harding & Muslimov, 2001). Soft X-ray pulsations at the known spin frequency of $271$ Hz were found by Ray et al. (2019) and further analyzed in Bogdanov et al. (2019a). The pulsed emission shows a broad and slightly asymmetric main pulse and a much fainter secondary interpulse (see Figure 1). This feature differs from those seen in NICER RMPs analyzed so far (see Bogdanov et al. 2019a for PSR J0030$+$0451 and PSR J0437--4715, and Wolff et al. 2021 for PSR J0740$+$6620). Similar to PSR J0740$+$6620 and PSR J0437--4715, PSR J1231--1411 is located in a binary system, allowing radio timing measurements to constrain the NS mass and observer inclination based on the relativistic Shapiro delay (Fonseca et al., 2021; Reardon et al., 2024). However, the obtained constraints in the preliminary radio timing results (available at the time of this work) are far less restrictive for PSR J1231--1411 (Cromartie et al. 2024, in preparation) compared to PSR J0740$+$6620 and PSR J0437--4715. In this work, we use these measurements as prior distributions and explore the robustness of NS parameter inference for PSR J1231--1411.
@@
Figure 1: The phase-folded PSR J1231--1411 event data for two rotational cycles (for clarity). The top panel shows the pulse profile summed over the channels. The total number of counts is given by the sum over all phase-channel pairs (over both cycles). For the modeling, all the event data are grouped into a single rotational cycle instead, and thus each phase-channel bin has twice the number of counts shown here.
Figure 2: XMM-Newton EPIC-pn camera PSR J1231--1411 count number spectrum as a function of nominal photon energies (blue-stepped curve). The measured rescaled background (red-stepped curve) and the background limits used in the analysis (purple dashed and dotted-dashed stepped curves) are also shown. The choice of the limits is explained in Section 3.4. The complete figure set (3 images), including also the spectra and backgrounds for EPIC-MOS1 and EPIC-MOS2 instruments is available in the online journal (HTML version).
@@ 2.1 NICER
[…]
@@
The timing model to assign pulse phases to photon measured times was developed by first fitting a timing model to the radio time of arrivals from the Nançay radio telescope (NRT). Details on the NRT and on pulsar observations with this instrument can be found in Guillemot et al. (2023). This determined the astrometric and binary parameters (including Shapiro delay) and a constant dispersion measure (DM). These were then held fixed while the frequency and frequency first derivative were fit to the Fermi-LAT gamma ray data, as done in FERMI-LAT Collaboration et al. (2022). Finally, the reference epoch was set to the values used in the Third Fermi-LAT Pulsar Catalog (3PC; Smith et al. 2023) so the absolute pulse phases will be consistent with that. This model provides pulse phases accurate to $<0.1\%$ of a pulse period over the entire duration of the NICER dataset.
@@
The NICER science tool nicerl3-spect was used to compute a pulse height spectrum, response matrix, and background estimate using the SCORPEON background model. [^19] https://heasarc.gsfc.nasa.gov/docs/nicer/analysis_threads/scorpeon-overview/ We used XSPEC (Arnaud, 1996) to fit the SCORPEON model parameters simultaneously with different source spectral models, including two blackbody components, one blackbody and one power-law component (with a power-law index fixed at $1.8$ as in Ransom et al. 2011), just a power-law component, one NS atmosphere component (nsatmos), and two NS atmosphere components. These were all modified by interstellar absorption (e.g., TBabs(BB + powerlaw)). For these data we found that, in addition to the standard free parameters in the SCORPEON model, we needed to free the normalization of the Ovii line at 0.57 keV (presumably originating from solar wind charge exchange). In prior work before the development of SCORPEON (Ray et al., 2019), this line was added as an ad hoc Gaussian line. We found that the obtained SCORPEON background estimates were notably dependent on the chosen source model, and therefore we predicted the background to range from the smallest estimate to the highest at each channel (after including also one standard deviation uncertainty for each model). This range was used only in the end to compare our inferred background to it (see Section 4.2).
@@
In the pulse profile analysis, we used the pulse invariant (PI) channel subset [30,150), corresponding to the nominal photon energy range [0.3, 1.5] keV, and 32 rotational phase bins. The data split over two rotational cycles are visualized in Figure 1.
@@ 2.2 XMM-Newton
The XMM-Newton EPIC-MOS (Turner et al. 2001; 1 and 2) and EPIC-pn (Strüder et al., 2001) data (ObsID 0605470201) used in this study were obtained on 2009 July 15, as part of a Fermi/LAT unidentified source observation campaign (Ransom et al., 2011). We utilize the Scientific Analysis System software (SAS) [^20] https://www.cosmos.esa.int/web/xmm-newton/sas to analyze and build our data products, which in turn are input into the inference software. The fields of the new unidentified LAT sources were observed using the XMM-Newton imaging mode, thus with insufficient time resolution to look for pulsations. Using the original raw data products from the EPIC-MOS and EPIC-pn observations, we extracted X-ray events from the now-known pulsar position. We set the PATTERN $\leq$ 12 for the EPIC-MOS1 and EPIC-MOS2 instruments and PATTERN $\leq$ 4 for the EPIC-pn along with FLAG=0 for all three instruments. This resulted in 28.8 ks of exposure from the EPIC-MOS instruments and 17.7 ks of exposure for the EPIC-pn instrument. For the EPIC-MOS1 and EPIC-MOS2 event data, we chose source+background extraction regions of $32\arcsec$ and source-free background extraction regions of $90\arcsec$. The extraction regions for source+background and background events for the EPIC-pn were $38\arcsec$ and $60\arcsec$, respectively. We filtered out times of higher-than-average particle noise from each event stream. The resulting final list of events are stored in FITS files that are then converted into ASCII files containing just the event times and PI value. However, no pulse phases are included in the XMM-Newton event files because the observations were made with a detector frame time that was too long to detect pulsations. Once the event filtering is selected, we build X-ray spectra using the SAS tool “evselect” and then use the SAS tools “rmfgen” and “arfgen” to build response matrices. The resulting response matrices are then turned into ASCII text files for ingesting into the inference code that is described below. We also implicitly make the assumption that the average pulsar X-ray properties have not changed from the epoch of the XMM-Newton observation to the epoch of the NICER observations. The assumption is justified since no evidence for long-term flux variability of PSR J1231--1411 was found in Bogdanov et al. (2019a, see their Section 7.3).
@@ 3 Modeling Procedure
We use the X-ray Pulse Simulation and Inference [^21] https://github.com/xpsi-group/xpsi (X-PSI) code, with version v2.1.1 for the inference runs and v2.2.7 for producing the figures (Riley et al., 2023). Complete information of each run, data products, posterior sample files, and all of the analysis files can be found in the Zenodo repository: doi:10.5281/zenodo.13358349 (Salmi et al., 2024). In the next sections, we summarize the modeling procedure and focus on how it differs from that used in previous work.
@@ 3.1 Effective-area Scaling Models
As in previous works (Riley et al., 2021; Choudhury et al., 2024a; Salmi et al., 2024; Vinciguerra et al., 2024), we capture the uncertainty in the effective area of both NICER and all XMM-Newton detectors by defining absolute energy-independent effective-area scaling factors. Each factor consists of a shared and telescope specific component (see Section 2.2. in Salmi et al. 2024), and, as in most recent works, we select the uncertainties in those to be 10% and 3%, correspondingly, leading to a 10.4% uncertainty in the full scaling factor. As before, we commonly assume all scaling factors to be identical between EPIC-pn, EPIC-MOS1, and EPIC-MOS2 (leaving then only the scaling factors $\alpha_{\mathrm{NICER}}$ and $\alpha_{\mathrm{XMM}}$ as free parameters). Since this may not be true, we have also performed a test run treating $\alpha_{\mathrm{XMM}}$ with $\alpha_{\mathrm{pn}}$, $\alpha_{\mathrm{MOS1}}$, and $\alpha_{\mathrm{MOS2}}$ as free independent parameters, although with all of them defined to have the same uncertainty (10.4% as mentioned above) and the same correlation with $\alpha_{\mathrm{NICER}}$ (see Section 4.1).
@@ 3.2 Pulse Profile Modeling Using X-PSI
As in previous NICER analyses, we use the “Oblate Schwarzschild” approximation for the spacetime and NS shape model to calculate the energy-resolved X-ray pulses (see e.g., Miller & Lamb, 1998; Nath et al., 2002; Poutanen & Gierliński, 2003; Cadeau et al., 2007; Morsink et al., 2007; Lo et al., 2013; AlGendy & Morsink, 2014; Bogdanov et al., 2019b; Watts, 2019). The spectrum and the beaming pattern of the surface radiation are in most cases assumed to follow those from a fully ionized hydrogen atmosphere model NSX (Ho & Lai, 2001), but we also performed a run assuming fully ionized helium NSX atmosphere (see Section 4.2). In addition, for one case we also modeled the NS surface outside the hot regions, assuming it emits blackbody radiation (for computational efficiency). This assumption is more realistic than having no surface emission, although more accurate still would be to use a numerical (and thus more expensive) partially ionized atmosphere model for the likely relatively cold surface. To account for the interstellar attenuation, we use the TBabs model (Wilms et al., 2000, updated in 2016).
To characterize the geometry (and number) of the hot regions, we apply two different models: Single Temperature-Unshared (ST-U) and Protruding Dual Temperature-Unshared (PDT-U). In the former model two hot regions are described by two single-temperature disjoint spherical caps with unshared parameters. The latter model is the most complex model applied with X-PSI so far, where each of the two hot regions is additionally described by two overlapping spherical caps with different temperatures (see Riley et al. 2019 and Figure 1 of Vinciguerra et al. 2023 for more details). Where the two overlap, emission of only one of them is accounted for, and that component is referred to as being “superseding” (the other being “ceding”).
For the pulse profile modeling resolution with X-PSI, we mostly follow the settings [^22] For more details see Section 2.3.1 in Vinciguerra et al. (2024) or X-PSI documentation at https://xpsi-group.github.io/xpsi/hotregion.html. used in the previous PSR J0740$+$6620 analyses (Riley et al., 2021; Salmi et al., 2022, 2024). For example, we calculate multiple images (photons reaching the observer from the same point at the NS surface with many deflection angles due to high gravity) up to third order in case they are visible. The initial deflection angle grid is again calculated for 512 different emission angles at each colatitude. However, to speed-up the calculation we reduced the number of model energies (from 128 to 64) and phases (from 64 to 32) for the more expensive PDT-U model. This change was tested and found to have no effects on the results for a ST-U run. In contrast, the resolution of the hot region cell mesh (sqrt_num_cells = 32, min_sqrt_num_cells = 10, max_sqrt_num_cells = 64) was not reduced to avoid loss of accuracy in the computed waveforms for extreme configurations, as found by Choudhury et al. (2024b).